### v1

# e-Profits: A Business-Aligned Evaluation Metric for Profit-Sensitive Customer Churn Prediction

Abstract
Retention campaigns in customer relationship management often rely on churn prediction models evaluated using traditional metrics such as AUC and F1-score. However, these metrics fail to reflect financial outcomes and may mislead strategic decisions. We introduce e-Profits, a novel business-aligned evaluation metric that quantifies model performance based on customer-specific value, retention probability, and intervention costs. Unlike existing profit-based metrics such as Expected Maximum Profit, which assume fixed population-level parameters, e-Profits uses Kaplan-Meier survival analysis to estimate personalised retention rates and supports granular, per-customer evaluation. We benchmark six classifiers across two telecom datasets (IBM Telco and Maven Telecom) and demonstrate that e-Profits reshapes model rankings compared to traditional metrics, revealing financial advantages in models previously overlooked by AUC or F1-score. The metric also enables segment-level insight into which models maximise return on investment for high-value customers. e-Profits is designed as an understandable, post hoc tool to support model evaluation in business contexts, particularly for marketing and analytics teams prioritising profit-driven decisions. All source code is available at: https://github.com/matifq/eprofits.
keywords:
Churn prediction , Machine learning & Artificial Intelligence , Profitable maximizing churn prediction , Profit based metrics
journal: Journal of Retailing and Consumer Services
[…]
1 Introduction
Customer churn, the end of a relationship between a customer and a business, presents a serious risk to a company’s sustainability. Whether customers choose to go to a competitor or stop using the service altogether, churn results in immediate revenue loss and a gradual decline in customer equity. Since attracting new customers is much more expensive typically five times more than keeping current ones [1, 2], many companies now focus on customer retention strategies enhanced by predictive analytics [3, 4]. Loyal customers may gradually acquire more goods and services over time and are associated with reduced marketing costs [5], and they spread positive sentiment which indirectly attract new customers. Therefore, higher customer retention rates result in increases in company profits .
Businesses increasingly rely on predictive models trained on behavioural and transactional data to address churn and identify customers at risk of leaving [6, 7]. Customer churn prediction (CPP) is typically framed as a binary classification task in the machine learning literature [8]. The goal is to identify would-be churners early enough to intervene, usually via targeted retention campaigns offering incentives. However, while predictive accuracy is important, most models are evaluated using traditional metrics such as F1-score or AUC, which do not reflect real-world business impact [9, 10]. This issue is especially problematic in profit-sensitive domains, where not all customers are equally valuable, and not all classification errors have the same cost [11, 12]. For instance, targeting low-value customers who were unlikely to churn wastes retention resources, while failing to identify high-value churners results in substantial lost revenue. Studies have shown that an inverse relationship may exist between customer lifetime value (CLV) and churn probability [13, 14], meaning that models can systematically overlook high-value customers optimised for statistical accuracy.
To bridge this gap, profit-based evaluation metrics have been proposed, including the Maximum Profit Criterion (MPC) and Expected Maximum Profit (EMP) [15, 16]. These metrics represent important progress but are often constrained by fixed assumptions, such as uniform CLV and constant offer acceptance probabilities, which rarely hold in real-world business contexts characterised by customer heterogeneity.
In this study, we introduce e-Profits, a business-aligned evaluation metric designed to assess churn models based on financial relevance rather than statistical precision. By incorporating customer-specific CLV, intervention cost, and retention probability, estimated through Kaplan-Meier survival analysis, e-Profits offers a granular, per-customer evaluation approach. We empirically evaluate six machine learning classifiers across two real-world telecom datasets (IBM Telco and Maven Telecom) and show that e-Profits reveals model rankings that diverge significantly from those based on traditional or existing profit-based metrics.
The remainder of the article is organised as follows. The next section 2 reviews the relevant literature. Following that, section 3 discusses the methodology and experimental setup. Sections 4 and 5 present the results and a discussion of their implications, respectively. Finally, the section 6 summarises the main findings of the work.
2 Related Works
Customer churn prediction has been extensively studied across domains such as telecommunications, banking, and subscription-based services [17]. Traditionally, churn is formulated as a supervised binary classification problem, with models trained to distinguish between customers who stay and those who leave. Early approaches employed logistic regression and decision trees, gradually evolving toward advanced techniques, including artificial neural networks [18], support vector machines [19], ensemble methods [20], and feature selection pipelines [21]. Despite these advances, the dominant focus remains on improving predictive performance measured through conventional metrics such as Accuracy, F1-score, and AUC [22].
These conventional metrics, however, are ill-suited for business settings where the cost of misclassification varies across customers. For example, AUC treats false positives and false negatives symmetrically [9], ignoring the financial impact of retention decisions. As Elkan noted in his foundational work on cost-sensitive learning [10], a classifier’s utility must be assessed based on domain-specific cost matrices rather than classification error alone. In churn prediction, misclassifying a high-value customer as a non-churner could lead to significant revenue loss, while misclassifying a low-value, loyal customer could result in wasted marketing resources.
Profit-based evaluation metrics have been introduced to address this gap. Verbeke et al. introduced the Maximum Profit Criterion (MPC) [15], which estimates profit based on classification outcomes. This was extended by Verbraken et al. through the Expected Maximum Profit (EMP) metric [16], which incorporates uncertainty in retention success and introduces the profit-maximising churn fraction, the optimal subset of customers to target. These metrics marked an important shift from accuracy-focused to business-aligned evaluation. Table 1 summarises the key characteristics of traditional and profit-aware evaluation metrics, highlighting the business relevance and flexibility offered by e-Profits compared to widely used alternatives.
Nevertheless, limitations persist. EMP assumes constant values for customer lifetime value (CLV), contact cost, and offer acceptance probability [23]. It models retention success using an unimodal Beta distribution with fixed hyperparameters, which may not capture the heterogeneity of real-world customers. Additionally, while some studies have explored embedding profit functions directly into model optimisation via profit-aware decision trees or evolutionary strategies [22, 24], these approaches often introduce complexity and reduce interpretability, limiting their uptake in business environments.
Some recent works have considered time-dependent churn risk using survival analysis models [25, 26]. These approaches explicitly demonstrate the value of incorporating tenure and retention time, but they seldom translate into actionable profit-based evaluation metrics.
Table 1: Comparison of churn evaluation metrics
Metric Type Business-Aware Fixed-Assumptions Individual-Level Optimisation Target
AUC Ranking No No Churn likelihood
F1-score Classification No No Class balance
EMP Profit-based Partial Yes No Profit-max subset
e-Profits Profit-based Yes No Yes Total and segment profit
While the academic literature contains several profit-aware metrics, their adoption in operational churn platforms remains rare. Surveys such as García et al. [27] and our own review [28] highlight that many practitioners still default to traditional metrics like AUC or lift, due to limitations in available tools and ease of understanding. This disconnect reinforces the need for practical and business-aligned evaluation metrics that can bridge the research-practice divide.
Conventional evaluation metrics such as Accuracy, F1-score, and AUC focus on statistical correctness but neglect the financial implications of churn prediction. Although profit-based metrics such as the EMP [16] represent an important step toward business-aligned evaluation, they often rely on population-level assumptions: fixed CLV, fixed intervention cost parameters (such as cost of contact and retention offer), and no integration of individualised retention rates. Such assumptions ignore the heterogeneity of real-world customer bases, where profit potential and intervention costs vary significantly across individuals.
Despite these advances, existing profit-aware approaches generally fail to capture customer heterogeneity. Most rely on fixed cost parameters or make strong assumptions about offer acceptance rates. By contrast, our proposed e-Profits is aimed at leveraging individual-level cost, revenue, and survival-based retention to provide a metric that is granular, understandable, and practically deployable, without retraining the model. e-Profits responds to these limitations by offering a modular evaluation metric that aligns directly with business impact, empowering practitioners to make financially sound model selection decisions.
3 Methodology and Experimental Setup
This section introduces our proposed evaluation metric, e-Profits, and explains how it aligns churn prediction with business objectives by incorporating customer-specific profitability and cost structure. We also introduce the details of a primary empirical study including the datasets selected, the classification techniques for churn prediction, and the comparative metrics for the purpose of validation.
3.1 Overview and Motivation
As established in the previous section, conventional evaluation metrics fail to capture the financial implications of churn misclassification, and existing profit-based alternatives often rely on strong population-level assumptions. In response, we introduce e-Profits, a modular evaluation metric that supports granular, business-aligned model comparison.
The e-Profits metric evaluates model performance by aggregating per-customer profit outcomes under realistic operational conditions. It incorporates:
  • Customer-specific lifetime value (CLV): Estimated using revenue and retention projections
  • Personalised retention probabilities: Derived via Kaplan-Meier survival analysis based on customer tenure
  • Individualised intervention costs: Including both contact and retention offer costs per customer
This design enables profit-based comparison of churn models without altering training pipelines, ensuring both understandability and practical deployment feasibility.
Model PredictionTP or FP?e-Profits = 0Cost of Intervention: - Cost of Offer - Cost of ContactTrue Positive?Calculate CLV using eq. 3e-Profits = CLV - Cost of offer- Cost of contacte-Profits = - Cost of offer - Cost of contactTotal e-Profits = $\sum$ e-ProfitsNoYesYesNo
Figure 1: Conceptual workflow of the e-Profits evaluation metric. This flow illustrates how model predictions interact with customer-level retention, CLV, and cost information to compute aggregate business impact. (To calculate CLV, we have used average and tenure-based retention rates (Using Kaplan-Meier))
3.2 e-Profits Formulation
We define e-Profits as the total estimated profit that would be realised if a retention campaign was executed based on the model’s churn predictions:
[…]
Where $\Pi_{i}$ is the estimated profit or loss associated with customer $i$, computed as:
$\Pi_{i}=\begin{cases}\text{CLV}_{i}-C_{\text{offer},i}-C_{\text{contact},i}&\text{if }y_{i}=1\text{ and }\hat{y}_{i}=1\quad\text{(True Positive)}\\ \\ -C_{\text{offer},i}-C_{\text{contact},i}&\text{if }y_{i}=0\text{ and }\hat{y}_{i}=1\quad\text{(False Positive)}\\ \\ 0&\text{otherwise (no retention action)}\end{cases}$ (2)
Where:
  • $y_{i}$ is the true churn label for customer $i$
  • $\hat{y}_{i}$ is the predicted label (1 = predicted to churn)
  • $\text{CLV}_{i}$ is the estimated customer lifetime value
  • $C_{\text{offer},i}$ and $C_{\text{contact},i}$ are the intervention costs for customer $i$
In this framework, retention actions are only initiated when $\hat{y}_{i}=1$ (i.e., the model predicts churn). True positives generate profit, while false positives incur unnecessary cost. No intervention, and hence no financial consequence, is applied to customers predicted as non-churners.
3.3 Customer Lifetime Value (CLV)
The CLV for each customer is calculated using the revenue, the estimated retention probability, and a fixed profit margin:
$\text{CLV}_{i}=\frac{R_{i}\times M}{1-\min(r_{i},0.995)}$ (3)
Where:
  • $R_{i}$ is the monthly revenue from customer $i$
  • $M$ is the profit margin
  • $r_{i}$ is the customer-specific retention probability
The minimum function ensures numerical stability by capping the retention probability at 0.995. This avoids division by values close to zero, which could lead to unrealistically large CLV estimates, especially in cases where long-tenured customers receive extremely high retention probabilities from survival models like Kaplan-Meier.
3.4 Retention Rate Estimation via Survival Analysis
To estimate the probability that a customer remains subscribed beyond a given tenure, we employ Kaplan-Meier survival analysis [25]. This method is non-parametric, meaning it does not assume any specific distribution of churn times, making it well-suited for business datasets where churn dynamics are irregular or unknown [29]. Compared to parametric or neural survival models, Kaplan-Meier offers a transparent and computationally efficient approach [30] that facilitates easy integration into evaluation workflows such as e-Profits.
We consider two forms of retention probability:
  • Average Retention Rate (ARR): A global retention probability calculated from the full dataset. This rate reflects the overall likelihood of a randomly selected customer staying subscribed and is computed using the aggregate churn durations in the dataset. In our case, this is based on each dataset’s customer records available at the time of data collection.
  • Tenure-based Retention Rate (TRR): A personalised retention estimate for each customer, derived from Kaplan-Meier survival curves conditioned on the customer’s current tenure (i.e., account length). This allows us to model retention more granularly, accounting for the fact that long-tenured customers are generally less likely to churn.
The CLV formula uses both ARR and TRR to compute expected profitability. To prevent unrealistically large CLV values due to extremely high retention estimates, we cap the retention rate at a maximum value of 0.995, ensuring numerical stability in the denominator, as discussed in Eq. 3.
3.5 Cost of Intervention
Intervention cost comprises two components: the cost of making contact with the customer and the cost of a retention offer. Both are proportional to the customer’s CLV.
$C_{\text{offer},i}=\text{CPO}\times\text{CLV}_{i}$ (4)
$C_{\text{contact},i}=\max(c_{0},c_{1}\times C_{\text{offer},i})$ (5)
Where:
  • CPO = Cost-per-offer ratio (e.g., 0.1)
  • $c_{0}$ = fixed minimum contact cost
  • $c_{1}$ = proportional multiplier for contact cost
This setup reflects realistic business conditions where retention costs scale with customer value. Complete workflow of e-profits is illustrated in figure 1.
3.6 Comparison with EMP
While the EMP metric also seeks to optimise profit, it differs from e-Profits in several key ways:
  • EMP assumes fixed CLV and global retention parameters [23]; e-Profits uses customer-specific inputs;
  • EMP models offer acceptance with a Beta distribution; e-Profits applies empirically estimated retention probabilities via survival analysis;
  • EMP is analytically optimised for a specific target fraction; e-Profits enables direct computation of total and top-segment profit without restrictive assumptions;
These features make e-Profits more understandable and suitable for operational decision-making in customer-centric businesses.
3.7 Implementation Guidance
e-Profits can be applied to any churn classifier without modifying the model architecture. It requires:
  1. 1.
    Customer-level predictions $\hat{y}_{i}$ and true labels $y_{i}$
  2. 2.
    Monthly revenue and tenure information to compute $\text{CLV}_{i}$
  3. 3.
    Retention rate estimation using Kaplan-Meier survival curves
  4. 4.
    Cost parameters for contact and offers
This modular design makes e-Profits readily integrable into standard evaluation pipelines and dashboards used by marketing and customer analytics teams.
3.8 Toy Example: Manual e-Profits Calculation
To illustrate how e-Profits can help evaluate profitability at the customer level, we present a toy example of five customers with known churn labels, predictions, and financial attributes. Only customers predicted as churners ($\hat{y}_{i}=1$) are considered for intervention, as per Eq. 3.
Table 2: Example customer data for e-Profits illustration
Customer $y_{i}$ (True) $\hat{y}_{i}$ (Pred) $R_{i}$ (Revenue) $r_{i}$ (Retention) Account Length
A 1 1 100 0.80 120
B 0 1 80 0.90 50
C 1 0 120 0.75 180
D 0 0 90 0.85 90
E 1 1 150 0.70 200
[…]
$\text{CLV}_{i}=\frac{R_{i}\cdot M}{1-min(r_{i},0.995)}$
The cost of intervention for each predicted churner is calculated using:
$C_{\text{offer},i}=\text{CPO}\cdot\text{CLV}_{i},\quad C_{\text{contact},i}=\max(c_{0},c_{1}\cdot C_{\text{offer},i})$
Customer A (TP):
$\text{CLV}_{A}=\frac{100\cdot 0.3}{1-0.80}=150,\quad C_{\text{offer}}=15,\quad C_{\text{contact}}=max(5,4.5)$
$\Pi_{A}=150-15-5.0=\mathbf{130}$
Customer B (FP):
$\text{CLV}_{B}=\frac{80\cdot 0.3}{1-0.90}=240,\quad C_{\text{offer}}=24,\quad C_{\text{contact}}=max(5,7.2)$
$\Pi_{B}=-24-7.2=\mathbf{-31.2}$
Customer C and D:
$\hat{y}_{C}=\hat{y}_{D}=0\Rightarrow\Pi_{C}=\Pi_{D}=0$
Customer E (TP):
$\text{CLV}_{E}=\frac{150\cdot 0.3}{1-0.70}=150,\quad C_{\text{offer}}=15,\quad C_{\text{contact}}=max(5,4.5)$
$\Pi_{E}=150-15-5=\mathbf{130}$
Total e-Profits:
$\sum\Pi_{i}=130+(-31.2)+0+0+130=\mathbf{228.8}$
This example illustrates how e-Profits accounts for both the financial value and the prediction correctness of each customer. While both Customers A and E contribute positively, Customer B, although flagged as a churner, incurs a financial loss due to being a false positive. By incorporating retention probability, revenue, and intervention cost, e-Profits provides a more financially meaningful evaluation than traditional classification metrics.
3.9 Datasets
We apply the e-Profits metric using two publicly available real-world churn datasets that meet the following criteria:
  • Actively used in recent churn prediction research
  • Publicly accessible for reproducibility
  • Moderate classification difficulty (accuracy or F1-score typically <85%)
1. IBM Telco Dataset [^2] https://www.kaggle.com/datasets/blastchar/telco-customer-churn/data This dataset contains 7,043 customer records from a telecommunications provider, with 19 predictive features including customer demographics, tenure, services subscribed, and payment methods. The churn rate is 26.5%.
2. Maven Telecom Dataset [^3] https://www.kaggle.com/datasets/johnp47/maven-churn-dataset This dataset includes 7,043 entries and 32 features. It presents a slightly different class distribution: 27% churners, 67% non-churners, and 6% recent joiners. We excluded recent joiners for binary classification. Both datasets contain no missing values and require minimal preprocessing. These datasets were selected to evaluate how e-Profits performs in settings with moderate class imbalance and diverse customer features.
3.10 Machine Learning Models
A prior systematic review of churn prediction techniques [28], tree-based ensemble data-driven learning techniques consistently demonstrated superior performance in predictive accuracy and robustness across various datasets. In particular, gradient-boosted techniques such as XGBoost, LightGBM, and CatBoost were frequently reported as top performers in learning accurate models, while Random Forest remained a strong baseline. Moreover, Explainable Boosting Machines were noted for offering a compelling balance between model accuracy and interoperability, an increasingly important consideration in decision-support applications [31]. In this study, we selected the following learning techniques for evaluation:
  • Random Forest (RF): Bagging-based ensemble proved to handle overfitting and high-dimensional data well [32].
  • Gradient Boosting Classifier (GBC): A boosting algorithm that sequentially reduces residual errors [33].
  • XGBoost (XGB): An optimised gradient boosting learning technique known for scalability and handling regularisation well[34].
  • LightGBM (LGBM): A leaf-wise boosting algorithm with histogram-based split search optimal for accelerated training [35].
  • CatBoost (CBC): a technique that can handle categorical variables natively and reduces prediction shift via ordered boosting [36].
  • Explainable Boosting Machine (EBM): A generalised additive model with competitive performance and enhanced interpretability [37].
All models were trained using a 5-fold cross-validation strategy given the datasets do not contain many records. Table 3 shows the hyperparameters and their ranges. This ensured fair performance comparison while accounting for model-specific tuning characteristics.
Table 3: Hyperparameter tuning grids for each classifier
Model Hyperparameters
RF n_estimators = [100, 200, 300]; max_depth = [None, 10, 20, 30]; min_samples_split = [2, 5, 10]; min_samples_leaf = [1, 2, 4]
XGB n_estimators = [100, 200, 300]; learning_rate = [0.01, 0.1, 0.2]; max_depth = [3, 4, 5]; colsample_bytree = [0.3, 0.7]
GBC n_estimators = [100, 200, 300]; learning_rate = [0.01, 0.1, 0.2]; max_depth = [3, 4, 5]
LGBM n_estimators = [100, 200]; learning_rate = [0.01, 0.1]; num_leaves = [31, 50]
CBC iterations = [100, 200]; learning_rate = [0.01, 0.1]; depth = [3, 5]
EBM learning_rate = [0.01, 0.05, 0.1]; max_bins = [128, 256]; interactions = [10, 20]; max_interaction_bins = [16, 32]
3.11 Evaluation Metrics
We benchmark e-Profits against both traditional and profit-oriented evaluation metrics, including:
  • Accuracy: Proportion of correctly classified instances. It is a simple metric, but it ignores class imbalance.
  • F1-score: Harmonic mean of precision and recall, appropriate for better handling imbalanced binary classification.
  • AUC (Area Under the ROC Curve): Measures ranking performance independent of decision threshold.
  • Top-Decile Lift: Measures how concentrated churners are in the top 10% of predicted probabilities.
  • Lift Index: Weighted index of churners across ranked deciles; values close to 1 indicate effective prioritisation.
  • Expected Maximum Profit (EMP): A profit-based metric assuming fixed CLV and offer acceptance probabilities [16].
  • e-Profits (ours): Our proposed metric that evaluates models based on customer-specific CLV, retention, and intervention cost.
Metrics are reported in two scenarios: the full population evaluation, which assesses the total profit across all predicted churners, and the top-segment evaluation, focusing on the profit generated from the top 20% of the most profitable customers based on predicted risk. This dual perspective allows us to understand both global model performance and targeted campaign efficiency [^4] Further mathematical formulations for traditional evaluation metrics can be found in our earlier review work [28].
4 Results
This section presents and analyses the performance of several machine learning classifiers for churn prediction using the two benchmark datasets: IBM Telco and Maven Telecom. Each classifier is evaluated using both traditional metrics (e.g., Accuracy, F1-score, AUC) and business-aligned metrics (e.g., Top-Decile Lift, Lift Index, EMP, and the proposed e-Profit. We aim to assess how model rankings and decision quality shift when profitability is considered alongside classification performance.
Results are organised in two parts. First, we report and discuss the results for the IBM Telco dataset, followed by the Maven dataset. We examine both full-population and top-segment outcomes for each dataset, focusing on how different tuning objectives (e.g., optimising for AUC vs. e-Profits) influence downstream profit impact.
4.1 IBM Dataset Analysis
We begin with the IBM Telco dataset, which contains customer tenure and service attributes relevant for churn prediction. Prior to model evaluation, we applied Kaplan-Meier survival analysis to estimate customer retention probabilities. The average retention rate across the IBM dataset is approximately 77%. Figure 2 provides a compact overview, combining the retention curve, lift index, and ROC curve for visual comparison.
(a) Kaplan-Meier estimated retention rate
(b) Lift Index chart using EBM
(c) ROC curves (tuned on e-Profits)
Figure 2: Overview of IBM dataset analysis: retention estimation, churn prioritisation, and classification performance.
To assess how well models prioritise high-risk churners, we computed the lift index for the best-performing model under default AUC tuning (EBM). The top-decile lift of 3.01 indicates that the top 10% of customers identified by the model are approximately three times more likely to churn than average, supporting targeted interventions.
Similarly, ROC curves under e-Profits tuning (Figure 2c) show that EBM delivers the best AUC, reaffirming its strong discrimination ability.
4.1.1 Classifier Performance across Metrics
Tables 4 to 8 report the evaluation outcomes for different models under five tuning objectives: AUC, Accuracy, F1-score, EMP, and e-Profits. For each setting, we report both full-population and top-segment e-Profits, as well as corresponding EMP, Lift, and classification metrics.
Table 4 reports results when classifiers are tuned using AUC. EBM outperforms others in classification metrics such as AUC, Accuracy, and F1-score. However, when ranked by top-segment e-Profits, CatBoost and LGBM exceed EBM. This indicates that AUC-optimised models may not always prioritise high-value customers effectively.
To enrich our analysis of model profitability, we report e-Profits using two variations of retention estimation: the global Average Retention Rate (ARR) and the customer-specific Tenure-based Retention Rate (TRR). Both are derived via Kaplan-Meier survival analysis, as described in the Methodology (see Section 3). This dual reporting allows us to examine how using personalised retention estimates affects model rankings under business-aligned metrics.
Table 4: Classifier performance on IBM dataset when tuned for AUC: Includes both traditional and profit-based evaluation metrics, including e-Profits (ARR/TRR) and EMP under full and top-segment scenarios.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
LGBM 56.25 78.80 69.82 10.65 2.9659 0.7981 25073 18952 224246 359836 33.35 10.65
RF 57.20 80.03 70.34 10.83 3.0008 0.8075 4532 3606 190975 335472 34.91 10.83
CBC 57.82 79.70 70.77 10.77 2.9833 0.8045 38184 29420 259429 390116 34.41 10.77
GBC 58.56 79.84 71.25 10.84 2.8787 0.8056 8910 7921 225756 374947 34.59 10.84
XGB 59.26 80.22 71.67 10.85 2.9659 0.8068 18365 13631 232821.29 378301 35.20 10.85
EBM 60.06 80.93 72.11 10.95 3.0182 0.8115 19292 15929 253196 399408 36.31 10.95
*CBC = CatBoost, e-prof = e-Profits, Acc. = Accuracy, TDL = Top Decile Lift, ARR = Average Retention Rate, TRR = Tenure Retention Rate
Table 5: Classifier performance on IBM dataset when tuned for Accuracy: Full-spectrum evaluation across classification, lift, and financial metrics, comparing e-Profits and EMP variants.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
LGBM 56.25 78.80 69.82 10.65 2.9659 0.7981 25073 18952 224246 359836 33.35 10.65
RF 57.20 80.03 0.7034 10.83 3.0008 0.8075 4532 3606 190975 335472 34.91 10.83
XGB 57.88 80.03 70.78 10.86 2.9659 0.8080 16144 12942 229247 372116 34.73 10.86
GBC 58.56 79.84 71.25 10.84 2.8787 0.8056 8910 7921 225756 374947 34.59 10.84
CBC 59.52 80.69 71.78 10.85 3.0008 0.8080 13276 10727 245988 390454 35.41 10.85
EBM 60.06 80.93 72.11 10.95 3.0182 0.8115 19292 15929 253196 399408 36.31 10.95
Table 6: Classifier performance on IBM dataset when tuned for F1-score: Financially-informed metrics reveal shifts in model rankings under e-Profits and EMP.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
LGBM 56.25 78.80 69.82 10.65 2.9659 0.7981 25073 18952 224246 359836 33.35 10.65
RF 57.20 80.03 70.34 10.83 3.0008 0.8075 4532 3606 190974 335472 34.91 10.83
CBC 57.82 79.70 70.77 10.77 2.9833 0.8045 38184 29420 259429 390116 34.41 10.77
GBC 58.56 79.84 71.25 10.84 2.8787 0.8056 8910 7921 225756 374947 34.59 10.84
XGB 59.26 80.22 71.67 10.85 2.9659 0.8068 18365 13631 232821 378301 35.20 10.85
EBM 60.06 80.93 72.11 10.95 3.0182 0.8115 19292 15929 253196 399408 36.31 10.95
Table 7: Classifier performance on IBM dataset when tuned for EMP: Traditional profit metric shows limitations when compared with e-Profits under personalised cost and value assumptions.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
XGB 19.78 75.44 55.28 10.91 2.7914 0.8059 0 0 13501 53261 34.41 10.91
GBC 32.60 76.72 59.16 10.65 2.6867 0.7951 0 0 28349 100746 32.34 10.65
CBC 44.58 77.99 63.76 10.76 2.7216 0.8031 0 0 76829 193024 33.62 10.76
RF 54.83 79.18 68.94 10.52 2.8961 0.7932 29209 20091 221862 343511 33.61 10.52
LGBM 56.25 78.80 69.82 10.65 2.9659 0.7981 25073 18952 224246 359836 33.35 10.65
EBM 58.60 80.41 71.20 11.00 2.9833 0.8099 8483 6959 219882 365990 35.20 11.00
Table 8: Classifier performance on IBM dataset when tuned directly for e-Profits: Demonstrates the financial benefit of model selection using customer-specific value, retention, and intervention costs.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
RF 54.10 78.56 68.51 10.42 2.8438 0.7918 20665 13917 221891 345815 32.05 10.42
LGBM 56.25 78.80 69.82 10.65 2.9659 0.7981 25073 18952 224246 359836 33.35 10.65
XGB 56.46 78.47 69.98 10.56 2.9135 0.7939 37805 27632 241516 372267 32.40 10.56
GBC 56.53 78.89 70.00 10.60 2.7565 0.7936 38196 29315 255144 382767 32.19 10.60
CBC 57.82 79.70 70.77 10.77 2.9833 0.8045 38184 29420 259429 390116 34.41 10.77
EBM 60.80 81.02 72.61 10.95 3.0182 0.8108 19445 15953 265068 413470 35.84 10.95
These results underscore that models optimised for ranking correctness may misallocate retention resources in real-world settings where profitability is paramount.
Tables 5 and 6 show model performance when tuned for Accuracy and F1-score, respectively. EBM again dominates classification performance, but the ranking for top-segment e-Profits shifts, with LGBM and CatBoost producing higher financial returns. This demonstrates that models selected using traditional metrics may misalign with business goals, especially in targeted retention campaigns.
These findings strengthen the case for evaluating classifiers by correctness and financial impact, especially in resource-constrained retention settings.
Table 7 provides results when models are tuned using EMP. While EMP reflects a profit orientation, its reliance on fixed cost and acceptance rate assumptions limits its alignment with individual-level profit signals. For instance, despite lower EMP, Random Forest delivers substantially better top-segment e-Profits than CatBoost. This inconsistency shows that even profit-based metrics like EMP can miss context-specific opportunities.
Notably, several models show zero profit in the top segment under EMP tuning, an artefact of global assumptions about CLV and offer response, not model capability.
Table 8 shows results when models are tuned directly on our proposed e-Profits metric. While EBM achieves the highest overall profit, GBC, CatBoost, and XGBoost yield superior outcomes for the top 20% most profitable customers. This is also illustrated by radar graph in figure 3. This divergence suggests that e-Profits is better suited for applications where ROI is maximised by targeting smaller, high-value segments.
By enabling per-customer evaluation using actual retention behaviour, cost, and value, e-Profits provides a practical basis for real-world model selection.
4.1.2 Key Insights
  • Traditional metrics (AUC, F1) fail to consistently rank models by business value.
  • EMP improves this alignment but remains sensitive to fixed parameter assumptions.
  • e-Profits offers a more faithful and granular evaluation by incorporating per-customer economics.
  • Classifier rankings change substantially when moving from AUC to e-Profits optimisation, highlighting the risk of metric misalignment in real-world deployment.
Figure 3: Radar graph for IBM dataset (Tuned on e-profits considering 20% profitable customers and tenure based retention)
4.2 Maven Dataset Analysis
We now turn to the Maven Telecom dataset, which features a richer set of features and a slightly different class distribution compared to IBM. We estimate customer retention probabilities using Kaplan-Meier analysis. The average retention rate in the Maven dataset is approximately 75%. Figure 4 summarises the retention estimates, lift behaviour, and classification performance.
(a) Kaplan-Meier estimated retention rate
(b) Lift Index chart using EBM
(c) ROC curves (tuned on e-Profits)
Figure 4: Overview of Maven dataset analysis: retention estimation, churn prioritisation, and classification performance.
The top-decile lift reaches 3.46, indicating effective risk segmentation. The ROC curve confirms that EBM achieves the highest AUC again, which is consistent with the findings of the IBM dataset.
4.2.1 Classifier Performance across Metrics
Classifier performance on the Maven dataset mirrors many trends observed on the IBM dataset. Full results are shown in Tables 9 to 13, comparing five tuning strategies and evaluating both global and top-segment e-Profits under ARR and TRR.
As with IBM, AUC-optimised models (Table 9) rank well in traditional metrics, but LightGBM and XGBoost surpass EBM in top-segment e-Profits. EMP-tuned models (Table 12) again fail to generalise well in high-value scenarios.
Most notably, Table 13 and Figure 5 demonstrates that when tuned directly on e-Profits, EBM yields the highest total profitability, while LightGBM captures the highest returns in the top segment. This reinforces e-Profits as a practical, fine-grained tool for selecting models that maximise business value.
Table 9: Classifier performance on Maven dataset when tuned for AUC: Includes traditional metrics, lift indicators, and business-aligned evaluation using e-Profits (ARR/TRR) and EMP under both full and top-segment scenarios.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
RF 73.71 86.65 80.18 13.09 3.4332 0.8521 6454 7533 28624 64093 47.18 13.09
GBC 75.00 86.75 81.45 13.25 3.4508 0.8589 7608 8272 30449 64715 48.084 13.25
CBC 75.19 86.95 81.49 13.44 3.4500 0.8616 7173 7910 30039 64919 48.03 13.44
LGBM 75.88 87.10 82.12 13.37 3.4500 0.8600 7874 8646 30971 65744 47.18 13.38
XGB 76.14 87.25 82.28 13.28 3.4700 0.8588 7704 8246 31370 66169 47.38 13.28
EBM 76.40 87.41 82.44 13.47 3.4684 0.8620 8009 8602 31412 66514 48.54 13.47
Table 10: Classifier performance on Maven dataset when tuned for Accuracy: Highlights classification quality alongside financial impact measured by e-Profits and EMP across customer segments.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
RF 73.26 86.60 79.78 13.14 3.4156 0.8532 6619 7636 27969 62905 47.35 13.14
GBC 75.00 86.75 81.45 13.25 3.4508 0.8589 7608 8272 30449 64715 48.08 13.25
CBC 75.19 86.95 81.49 13.44 3.4508 0.8616 7173 7910 30039 64919 48.03 13.44
XGB 75.41 87.00 81.68 13.41 3.4684 0.8626 7125 7636 30495 64987 48.54 13.41
LGBM 75.88 87.10 82.12 13.37 3.4508 0.8598 7874 8646 30971 65744 47.18 13.37
EBM 76.40 87.41 82.44 13.47 3.4683 0.8619 8009 8602 31412 66514 48.54 13.47
Table 11: Classifier performance on Maven dataset when tuned for F1-score: Shows how financial metrics like e-Profits shift model preferences compared to standard evaluation.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
RF 73.71 86.65 80.18 13.09 3.4332 0.8521 64544 7533 28624 64093 47.18 13.09
GBC 75.00 86.75 81.45 13.25 3.4508 0.8589 7608 8272 30449 64715 48.08 13.25
CBC 75.19 86.95 81.49 13.44 3.4508 0.8616 7173 7910 30039 64919 48.03 13.44
XGB 75.86 87.25 81.97 13.37 3.4684 0.8604 7366 7885 30387 64963 48.25 13.37
LGBM 75.88 87.10 82.12 13.37 3.4508 0.8598 7909 8646 30971 65744 47.18 13.37
EBM 76.40 87.41 82.44 13.47 3.4684 0.8619 8008 8602 31412 66514 48.54 13.47
Table 12: Classifier performance on Maven dataset when tuned for EMP: EMP’s reliance on fixed assumptions is contrasted with e-Profits’ granular evaluation across ARR and TRR.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
XGB 47.94 80.22 65.75 12.79 3.4508 0.8453 0.00 0.00 9311 34366 44.97 12.79
GBC 58.48 82.90 70.66 12.77 3.4156 0.8460 1974 2145 15567 43252 46.16 12.77
CBC 66.59 84.37 75.35 12.76 3.3980 0.8475 3262 3763 218045 54852 45.48 12.76
RF 73.41 86.44 80.04 12.91 3.4156 0.8518 6567 7601 28318 63747 47.18 12.91
LGBM 75.88 87.10 82.12 13.37 3.4508 0.8598 7909 8646 30971 65744.38 47.18 13.37
EBM 76.40 87.41 82.44 13.47 3.4684 0.8619 8008 8602 31412 66514 48.54 13.47
Table 13: Classifier performance on Maven dataset when tuned directly for e-Profits: Demonstrates effectiveness of personalised financial evaluation using e-Profits under different retention assumptions.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
RF 73.71 86.65 80.18 13.09 3.4332 0.8521 6454 7533 28624 64093 47.18 13.09
GBC 75.00 86.75 81.45 13.25 3.4508 0.8589 7608 8272 30449 64715 48.08 13.25
CBC 75.19 86.95 81.49 13.44 3.4508 0.8616 7173 7910 30040 64919 48.03 13.44
LGBM 75.88 87.10 82.12 13.37 3.4508 0.8598 7909 8646 30971 65744 47.18 13.37
XGB 76.14 87.25 82.28 13.28 3.4684 0.8588 7704 8246 31370 66169 47.38 13.28
EBM 76.40 87.41 82.44 13.47 3.4684 0.8619 8008 8602 31412 66514 48.54 13.47
4.2.2 Key Insights
  • Results from the Maven dataset reinforce the limitations of traditional metrics for model selection in profit-sensitive scenarios.
  • EMP tuning, although profit-aware, is constrained by its fixed assumptions and fails to highlight top-performing models in high-value segments.
  • e-Profits again provides a clearer differentiation, especially for top-segment profitability, helping identify models like LightGBM that may otherwise be overlooked.
  • The consistency of findings across datasets highlights e-Profits as a reliable, business-aligned evaluation framework.
Figure 5: Radar graph for Maven dataset (Tuned on e-profits considering 20% profitable customers and tenure based retention)
5 Discussion
Our findings across the IBM and Maven datasets reveal consistent patterns that underscore the practical value of using e-Profits as an evaluation metric for churn prediction models. Traditional performance metrics such as AUC and F1-score did not consistently identify the most financially effective models. For instance, classifiers like LGBM and CatBoost, which did not top the rankings under accuracy or AUC, emerged as highly profitable when evaluated using e-Profits. This reordering highlights a critical discrepancy between predictive correctness and financial value, underscoring the limitations of relying solely on statistical metrics for real-world business decisions.
Moreover, e-Profits proved especially effective in identifying models that excel within high-value customer segments. In both datasets, CatBoost and GBC generated the highest profits when analysis was focused on the top 20% of customers sorted by predicted churn risk and estimated profitability. These models were not necessarily the best according to traditional metrics, reinforcing the importance of adopting evaluation criteria that prioritise return on investment (ROI) over aggregate accuracy. The ability to tailor model evaluation towards strategic, high-impact interventions makes e-Profits a valuable tool for resource-constrained retention campaigns.
A key insight from our analysis is the limited utility of commonly used lift metrics in reflecting actual business outcomes. While lift remains a popular heuristic in industry, it often fails to account for the heterogeneity in customer value and the variable cost of retention. For example, Table 7 illustrates scenarios where models with strong lift scores delivered zero e-Profits, indicating that churn likelihood alone is an insufficient criterion for profitable targeting. By combining churn risk with revenue potential and cost structure, e-Profits offers a more holistic and financially meaningful basis for model selection.
In addition to these strategic insights, we also note several practical considerations in the implementation of e-Profits. First, accurate estimation of customer lifetime value and retention probabilities requires access to detailed data, including monthly revenue, tenure, and timestamped churn labels. Organisations may need to enrich existing customer datasets to support such granular modelling. Second, e-Profits is sensitive to cost parameters such as contact cost and the cost of providing retention offers. While our experiments employed domain-informed heuristics for offer costs (e.g., proportions of CLV informed by CRM practices), businesses may need to calibrate these parameters based on internal financial policies and strategic goals. Third, our current approach assumes a binary intervention model, deciding whether or not to contact a customer based on predicted churn. Future extensions could consider multi-tiered or continuous intervention strategies to capture more complex business realities.
We also justified the use of Kaplan-Meier survival analysis for retention rate estimation, citing its simplicity, non-parametric nature, and understandability. These qualities make it suitable for business contexts where model transparency and ease of implementation are key. Although more complex methods such as DeepSurv [26] and DRSA [38] may offer improved predictive accuracy, their integration typically requires changes to model training workflows and increased computational overhead. For the purposes of post hoc model evaluation, Kaplan-Meier offers a robust and practical trade-off, delivering reliable customer-level retention estimates without excessive complexity.
In summary, e-Profits extends beyond predictive performance to deliver insights directly aligned with financial decision-making. Its application enables marketing and analytics teams to select models that maximise business impact rather than statistical precision alone, positioning it as a valuable addition to the model evaluation toolkit in customer retention contexts.
6 Conclusion
Traditional churn prediction models are often assessed using statistical performance metrics such as Accuracy, AUC, and F1-score. While these metrics serve well in quantifying classification performance, they frequently fail to capture the real financial consequences of misclassification, particularly in profit-sensitive applications like customer retention.
In this study, we introduced e-Profits, a business-aligned evaluation metric that integrates customer-specific lifetime value (CLV), intervention costs, and retention probabilities derived from Kaplan-Meier survival analysis. Unlike traditional metrics and even profit-based alternatives such as EMP, e-Profits evaluates models based on individualised cost-revenue dynamics, facilitating better alignment between predictive analytics and financial outcomes. Our empirical evaluation on two benchmark datasets demonstrated that e-Profits significantly reshapes classifier rankings. Models that were suboptimal under Accuracy or AUC often emerged as the most financially effective under e-Profits, particularly when focusing on the top 20% of customers by predicted profitability. This shift underscores the potential for misaligned decision-making if only conventional metrics are used. The metric’s post hoc nature and lightweight implementation make e-Profits practical for real-world deployment. It requires no retraining and can be integrated easily into existing evaluation workflows, making it accessible to business analysts and operational stakeholders alike. Moreover, its understandable structure enhances communication between technical teams and decision-makers.
Future directions for e-Profits include embedding it as an objective function during model training and experimenting with advanced survival models for improved retention prediction. Overall, e-Profits introduces a financially meaningful evaluation paradigm that empowers practitioners to make data-driven decisions rooted in return on investment.
CRediT authorship contribution statement
Awais Manzoor: Conceptualization, Data curation, Methodology and Experimental setup, Software, Formal analysis, Investigation, Validation, Writing – original draft, Writing – review & editing M. Atif Qureshi: Conceptualization, Data curation, Methodology and Experimental setup, Software, Formal analysis, Investigation, Validation, Writing – review & editing, Supervision Etain Kidney: Supervision, Writing – review & editing Luca Longo: Supervision, Writing – review & editing
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This publication has emanated from research conducted with the support of Research Ireland under award no. GOIPG/2021/1354, and with the financial support of Research Ireland under grant no. 13/RC/2106_P2 at the ADAPT Research Centre at Technological University Dublin.